Two cars begin driving from the same point in the opposite directions. The first car travels at an average rate of 58 miles per hour. The second car drives 67 miles per hour. How many minutes will it take for the cars to be 100 miles apart?
I started this but I don't think it's right:
100/58+60=t
Your help is greatly appreciated!!

Answer by scott8148(6628)

they are moving apart at 125 mph (67 + 58)

d = r * t ___ 100 = 125 * t ___ the t will be in hours, so convert it to minutes